OpenStudy (anonymous):

Find the slope m of the tangent to the curve y = 5/sqrt(x) at the point where x = a > 0.

zepdrix (zepdrix):

The slope of the line tangent to \(\Large\rm y(x)\) at x=a will be given by \(\Large\rm y'(a)\).

zepdrix (zepdrix):

Hmm, so we need to take a derivative. Understand how to do that?

OpenStudy (anonymous):

the derivative is \[5\div x ^{2}\]

zepdrix (zepdrix):

Hmm that looks a little bit off :O

zepdrix (zepdrix):

\[\Large\rm y=5x^{-1}\]We should be getting a negative showing up somewhere when we differentiate, yes?

OpenStudy (anonymous):

\[-5\div 2a ^{5/2}\]

zepdrix (zepdrix):

Hmm that looks kinda strange D:

zepdrix (zepdrix):

Was this the function we started out with?\[\Large\rm y=\frac{5}{x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh no im sorry it was 5\[5\div \sqrt{x}\]

zepdrix (zepdrix):

\[\Large\rm y=\frac{5}{\sqrt{x}}\]Let's write our x with a rational exponent:\[\Large\rm y=\frac{5}{x^{1/2}}\]And bring it up to the numerator:\[\Large\rm y=5x^{-1/2}\]Now we can easily apply our power rule to find the derivative.

zepdrix (zepdrix):

\[\Large\rm y'=-\frac{1}{2}\cdot 5x^{(-\frac{1}{2}-1)}\]Power rule, yes? :)

OpenStudy (anonymous):

y'= \[-1\div2 \times5x ^-3\div2\]

zepdrix (zepdrix):

Mmm good good good :)

zepdrix (zepdrix):

So your solution was very close. Looks like you got a -5/2 the first time by mistake instead of -3/2 for the exponent.

OpenStudy (anonymous):

thank you